# On the Value of Cross-Modal Misalignment in Multimodal Representation Learning

### v6

@@ 4.2 Insights into Cross-Modal Misalignment for Practice
Assume that Asms.˜3.1 and 3.2 hold. Let the selection bias be $\theta=2^{n_{s}}-1$ and the perturbation bias be $\rho=1$, such that the full set of semantic variables $\mathbb{I}_{\mathbf{s}}$ is selected, and the perturbable semantic subset is trivial, i.e., $\mathbb{I}_{\rho}=\emptyset$. Then, all semantic variables $\mathbf{s}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n_{s}}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n_{s}}$, when minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$.
@@
Assume that Asms.˜3.1 and 3.2 hold. Consider an OOD setting in which a subset of semantic variables, $\mathbb{I}_{inv}\subset\mathbb{I}_{\mathbf{s}}$, remains invariant between training and testing environments, while the remaining semantic variables, $\mathbb{I}_{var}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{inv}$, undergo distribution shifts. If the union of omitted and perturbable semantic variables under selection bias $\theta$ and perturbation bias $\rho$ coincides with the environment-sensitive subset, i.e., $\mathbb{I}_{var}=\mathbb{I}_{\theta}^{c}\cup\mathbb{I}_{\rho},$ then the invariant semantic variables $\mathbf{s}_{\mathbb{I}_{inv}}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{|\mathbb{I}_{inv}|}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{|\mathbb{I}_{inv}|},$ by minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$.
@@ 5.1 Numerical Simulation
We simulate data following the generative process in §˜3. Semantic variables $\mathbf{s}\sim\mathcal{N}(0,\Sigma_{\mathbf{s}})$ (dim. 10) and modality-specific variables $\mathbf{m}_{x}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{x}}),\mathbf{m}_{t}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{t}})$ (dim. 5) encode potential causal dependencies via their respective covariances $\Sigma_{(\cdot)}$. To study cross-modal misalignment, we independently vary: (i) Selection bias: controlling $\mathbb{I}_{\theta}=\{1\},\dots,[10]$, and (ii) Perturbation bias: defining $\mathbb{I}_{\rho}=\emptyset,\dots,[9]$, with Gaussian noise added to each $i\in\mathbb{I}_{\rho}$ with probability 0.75. In isolation, we fix $\mathbb{I}_{\rho}=\emptyset$ when varying $\mathbb{I}_{\theta}$, and $\mathbb{I}_{\theta}=[10]$ when varying $\mathbb{I}_{\rho}$. Image and text data are generated using randomly initialized invertible MLPs $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle\theta}}$. We train two modality-specific MLP encoders for 100,000 steps using the $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ loss in Eq.˜2, with embedding dimension matching the number of unbiased semantic components. See §˜D.1 for further details.
@@ Appendix A Notation and Terminology
$\mathbf{g}_{x}$ Differomorphic generative mapping for images: $\mathcal{S}\times\mathcal{M}_{x}\to\mathcal{X}$
$\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ Differomorphic generative mapping for text under selection $\theta$: $\mathcal{S}_{\mathbb{I}_{\theta}}\times\mathcal{M}_{t}\to\mathcal{T}^{\scriptscriptstyle(\theta)}$
$\mathcal{G}_{t}$ A diffeomorphism function class where $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ is selected from
$\mathbf{f}_{x}$ Image encoder: $\mathcal{X}\to(0,1)^{n}$ with specified $n$
$\mathbf{f}_{t}$ Text encoder: $\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ with specified $n$
@@
$\mathcal{L}_{\text{\tiny MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Symmetric InfoNCE loss for MMCL (Eq.˜1)
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Alignment and entropy maximization loss (Eq.˜2)
@@ Appendix B Related Work
Identifiability analysis addresses the fundamental question of whether the learning process can uniquely recover the latent generative structure or distribution underlying the observed data. This problem has been extensively studied in the context of nonlinear independent component analysis (ICA) [23, 46, 28, 63] and causal representation learning (CRL) [41, 43, 80, 77, 44]. In either setting, full identifiability—typically up to permutation—is rarely achievable in practice without strong assumptions, such as access to interventional data. Consequently, recent works have focused on partial identifiability [20, 31, 18] or relaxed equivalence classes, such as identifiability up to linear transformations [83, 45] or up to group-wise/block-wise indeterminacy [68, 78], which can offer sufficient guarantees for specific tasks or applications. In the context of multimodal representation learning, several recent studies have explored identifiability results [48, 13, 45], but largely neglect the presence of cross-modal misalignment. In contrast, our work explicitly models misalignment and adopts a block-identifiability definition to characterize the extent to which semantic factors can be recovered up to an invertible mapping.
@@
Invariant representation learning (IRL) seeks to learn representations that remain robust under distributional shifts between environments [2, 15, 18], particularly in settings where empirical risk minimization (ERM) [66] fails to generalize out of distribution. In the absence of such challenge, ERM is sufficient for in-distribution prediction, rendering the objective of IRL unnecessary. From a causal perspective, learning invariant representations—or more ambitiously, autonomous mechanisms [60]—requires observing sufficient variations in the latent factors underlying data [41]. Such variability can be introduced through interventional data [39, 33], exchangeability assumptions [56], or the use of auxiliary variables such as domain indices [80, 42]. However, direct interventions on latent variables are typically infeasible in real-world observational data, and auxiliary variable methods often require access to a large number of diverse environments to ensure identifiability—an assumption that is often challenging to satisfy in practice. Our work sheds light on an alternative approach: by leveraging the inherent flexibility of text supervision, we show that manipulating biases, specifically through selective omission or semantic perturbation in text, can serve as a controllable proxy for environmental variation.
@@ C.1 Lemmas
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\left[\|\mathbf{f}_{x}(\mathbf{x})-\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\right]-\frac{1}{2}\Bigl{(}H\bigl{(}\mathbf{f}_{x}(\mathbf{x})\bigr{)}+H\bigl{(}\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr{)}\Bigr{)},$ (6)
@@
$\mathbf{f}_{x}^{*}=\mathbf{d}\circ(\mathbf{g}^{-1}_{x})_{\mathbb{I}^{c}_{\rho}}:\mathcal{X}\to(0,1)^{n},$ (7)
$\mathbf{f}_{t}^{*}=\mathbf{d}\circ(\mathbf{g}^{-1}_{t^{\scriptscriptstyle(\theta)}})_{\mathbb{I}^{c}_{\rho}}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n},$ (8)
@@
  • $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ denote the true underlying generative mappings for images and paired text, respectively, as described in §˜3.
@@
We prove that the candidate functions $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$ in Equations˜7 and 8 yield $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0$. Substituting these candidate functions into the loss in Eq.˜6, we have
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\Bigl{[}\|\mathbf{f}_{x}^{*}(\mathbf{x})-\mathbf{f}_{t}^{*}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\Bigr{]}$
$\quad-\frac{1}{2}\Bigl{(}H\bigl{(}\mathbf{f}_{x}^{*}(\mathbf{x})\bigr{)}+H\bigl{(}\mathbf{f}_{t}^{*}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr{)}\Bigr{)}.$
By the invertibility of the generative processes $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ (see §˜3), we may change variables to express the expectation over the latent variables:
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\sim p_{\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}}}\Bigl{[}\|\mathbf{d}(\mathbf{s}_{\mathbb{I}^{c}_{\rho}})-\mathbf{d}(\tilde{\mathbf{s}}_{\mathbb{I}^{c}_{\rho}})\|_{2}\Bigr{]}$
@@
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0.$
Therefore, the global minimum of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ is achieved at 0 by the given function pairs $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$, completing the proof. ∎
@@
Let $\mathbf{h}:\mathcal{U}\to\mathcal{V}$ be a smooth map between simply connected, open $\mathcal{C}^{1}$ manifolds $\mathcal{U},\mathcal{V}\subseteq\mathbb{R}^{n}$. Suppose that $\mathbf{u}$ is a random variable taking values in $\mathcal{U}$ with a smooth probability density that is strictly positive a.e.. If the pushforward $\mathbf{v}=\mathbf{h}(\mathbf{u})$ is uniformly distributed on $\mathcal{V}$, then $\mathbf{h}$ is a global diffeomorphism; in other words, $\mathbf{h}$ is bijective and depends on every component of $\mathbf{u}$. [^8] We do not claim originality for this result due to its fundamental nature in topology and measure theory; rather, we detail it here as a tool for our subsequent arguments.
@@
Let $p_{\mathbf{u}}:\mathcal{U}\to\mathbb{R}$ and $p_{\mathbf{v}}:\mathcal{V}\to\mathbb{R}$ denote the probability density functions of $\mathbf{u}$ and $\mathbf{v}$, respectively. Since $\mathbf{v}=\mathbf{h}(\mathbf{u})$, the change-of-variables formula (refer to, e.g., [5]) yields
$p_{\mathbf{v}}(\mathbf{v})=p_{\mathbf{u}}(\mathbf{u})\cdot\Bigl{|}\det J(\mathbf{h})(\mathbf{u})\Bigr{|}^{-1},$
where $\mathbf{u}$ is any preimage of $\mathbf{v}$ under $\mathbf{h}$. By assumption, $\mathbf{v}$ is uniformly distributed on $\mathcal{V}$; that is, there exists a constant $C>0$ such that
@@
Thus, for any $\mathbf{u}$ with $\mathbf{v}=\mathbf{h}(\mathbf{u})$ we obtain
$\Bigl{|}\det J(\mathbf{h})(\mathbf{u})\Bigr{|}^{-1}=\frac{C}{p_{\mathbf{u}}(\mathbf{u})}.$
@@
$\Bigl{|}\det J(\mathbf{h})(\mathbf{u})\Bigr{|}^{-1}>0\quad\text{a.s.}\quad\forall\,\mathbf{u}\sim p_{\mathbf{u}},$
or equivalently, $\det J(\mathbf{h})(\mathbf{u})\neq 0$ a.e.. By the Inverse Function Theorem (see, e.g., [58]), this implies that $\mathbf{h}$ is a local diffeomorphism.
Moreover, since $\mathcal{U}$ and $\mathcal{V}$ are simply connected, open $\mathcal{C}^{1}$ manifolds, standard covering space theory (refer to, e.g., the discussion around Theorem 1.38 in [21]) implies that $\mathbf{h}$ is a covering map. The uniformity of $p_{\mathbf{v}}$ forces $\mathbf{h}$ to be surjective (otherwise, some points in $\mathcal{V}$ would have zero density, contradicting uniformity). Since any covering map from a simply connected space is trivial, $\mathbf{h}$ must be equivalent to the identity covering. In other words, $\mathbf{h}$ is a homeomorphism onto $\mathcal{V}$ and hence both injective and surjective (i.e., a bijection).
Finally, the fact that the Jacobian determinant is nonzero a.e. guarantees that $\mathbf{h}$ depends on all components of $\mathbf{u}$; if any component were omitted, the rank of the Jacobian would drop, contradicting non-singularity. Furthermore, by the Global Inverse Function Theorem (refer to, e.g., [59]), the inverse of $\mathbf{h}$ is smooth.
In summary, $\mathbf{h}$ is a global diffeomorphism from $\mathcal{U}$ onto $\mathcal{V}$. Consequently, it preserves all information of $\mathbf{u}$: every variation in $\mathbf{u}$ is reflected in $\mathbf{v}=\mathbf{h}(\mathbf{u})$, and $\mathbf{u}$ can be uniquely and smoothly recovered from $\mathbf{v}$. This completes the proof. ∎
@@
  1. 1.
    First, we show that the objective function $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ achieves a global minimum value of 0. At this minimum, any pair of smooth functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ satisfying this condition must exhibit invariance across modalities. This invariance condition ensures that the learned image representations and text representations must align across all positive $\mathbf{x}$ and $\mathbf{t}^{\scriptscriptstyle(\theta)}$ pairs.
  2. 2.
@@
Step 1 (Global Minimum and Invariance Condition). Let $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ be any smooth functions attaining the global minimum. Define the smooth mappings:
$\mathbf{h}_{x}=\mathbf{f}_{x}\circ\mathbf{g}_{x},\quad\mathbf{h}_{t}=\mathbf{f}_{t}\circ\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}.$
@@
$\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x}}\,p_{\mathbf{t}^{\scriptscriptstyle(\theta)}|\mathbf{x}}}\left[\|\mathbf{f}_{x}(\mathbf{x})-\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\right]=\mathop{\mathbb{E}}_{(\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\sim p_{\mathbf{z}_{x}}p_{\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}|\mathbf{z}_{x}}}\left[\|\mathbf{h}_{x}(\mathbf{z}_{x})-\mathbf{h}_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\|_{2}\right]=0,$ (9)
$H\bigl{(}\mathbf{f}_{x}(\mathbf{x})\bigr{)}=H\bigl{(}\mathbf{h}_{x}(\mathbf{z}_{x})\bigr{)}=0,$ (10)
$H\bigl{(}\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr{)}=H\bigl{(}\mathbf{h}_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\bigr{)}=0.$ (11)

### v7

@@ 4.2 Insights into Cross-Modal Misalignment for Practice
Assume that Asms.˜3.1 and 3.2 hold. Let the selection bias be $\theta=2^{n_{s}}-1$ and the perturbation bias be $\rho=1$, such that the full set of semantic variables $\mathbb{I}_{\mathbf{s}}$ is selected, and the perturbable semantic subset is trivial, i.e., $\mathbb{I}_{\rho}=\emptyset$. Then, all semantic variables $\mathbf{s}$ are block-identified via smooth functions $f_{x}:\mathcal{X}\to(0,1)^{n_{s}}$ and $f_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n_{s}}$, when minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$.
@@
Assume that Asms.˜3.1 and 3.2 hold. Consider an OOD setting in which a subset of semantic variables, $\mathbb{I}_{inv}\subset\mathbb{I}_{\mathbf{s}}$, remains invariant between training and testing environments, while the remaining semantic variables, $\mathbb{I}_{var}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{inv}$, undergo distribution shifts. If the union of omitted and perturbable semantic variables under selection bias $\theta$ and perturbation bias $\rho$ coincides with the environment-sensitive subset, i.e., $\mathbb{I}_{var}=\mathbb{I}_{\theta}^{c}\cup\mathbb{I}_{\rho},$ then the invariant semantic variables $\mathbf{s}_{\mathbb{I}_{inv}}$ are block-identified via smooth functions $f_{x}:\mathcal{X}\to(0,1)^{|\mathbb{I}_{inv}|}$ and $f_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{|\mathbb{I}_{inv}|},$ by minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$.
@@ 5.1 Numerical Simulation
We simulate data following the generative process in §˜3. Semantic variables $\mathbf{s}\sim\mathcal{N}(0,\Sigma_{\mathbf{s}})$ (dim. 10) and modality-specific variables $\mathbf{m}_{x}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{x}}),\mathbf{m}_{t}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{t}})$ (dim. 5) encode potential causal dependencies via their respective covariances $\Sigma_{(\cdot)}$. To study cross-modal misalignment, we independently vary: (i) Selection bias: controlling $\mathbb{I}_{\theta}=\{1\},\dots,[10]$, and (ii) Perturbation bias: defining $\mathbb{I}_{\rho}=\emptyset,\dots,[9]$, with Gaussian noise added to each $i\in\mathbb{I}_{\rho}$ with probability 0.75. In isolation, we fix $\mathbb{I}_{\rho}=\emptyset$ when varying $\mathbb{I}_{\theta}$, and $\mathbb{I}_{\theta}=[10]$ when varying $\mathbb{I}_{\rho}$. Image and text data are generated using randomly initialized invertible MLPs $g_{x}$ and $g_{t^{\scriptscriptstyle\theta}}$. We train two modality-specific MLP encoders for 100,000 steps using the $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ loss in Eq.˜2, with embedding dimension matching the number of unbiased semantic components. See §˜D.1 for further details.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
The authors gratefully acknowledge support from the Responsible AI Research (RAIR) Centre (J. Q. Shi and Z. Zhang), the Centre for Augmented Reasoning (CAR) (Y. Liu and E. Gao), and the Commonwealth Bank of Australia through the CommonBank Centre for Foundational AI Research (A. van den Hengel). This support was essential to the completion of the research presented in this publication. The authors also thank the anonymous reviewers for their constructive feedback.
@@ Appendix A Notation and Terminology
$g_{x}$ Differomorphic generative mapping for images: $\mathcal{S}\times\mathcal{M}_{x}\to\mathcal{X}$
$g_{t^{\scriptscriptstyle(\theta)}}$ Differomorphic generative mapping for text under selection $\theta$: $\mathcal{S}_{\mathbb{I}_{\theta}}\times\mathcal{M}_{t}\to\mathcal{T}^{\scriptscriptstyle(\theta)}$
$\mathcal{G}_{t}$ A diffeomorphism function class where $g_{t^{\scriptscriptstyle(\theta)}}$ is selected from
$f_{x}$ Image encoder: $\mathcal{X}\to(0,1)^{n}$ with specified $n$
$f_{t}$ Text encoder: $\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ with specified $n$
@@
$\mathcal{L}_{\text{\tiny MMCL}}(f_{x},f_{t})$ Symmetric InfoNCE loss for MMCL (Eq.˜1)
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(f_{x},f_{t})$ Alignment and entropy maximization loss (Eq.˜2)
@@ Appendix B Related Work
Identifiability analysis addresses the fundamental question of whether the learning process can uniquely recover the latent generative structure or distribution underlying the observed data. This problem has been extensively studied in the context of nonlinear independent component analysis (ICA) [23, 46, 28, 63] and causal representation learning (CRL) [41, 43, 80, 77, 44]. In either setting, full identifiability, typically up to permutation, is rarely achievable in practice without strong assumptions, such as access to interventional data. Consequently, recent works have focused on partial identifiability [20, 31, 18] or relaxed equivalence classes, such as identifiability up to linear transformations [83, 45] or up to group-wise/block-wise indeterminacy [68, 78], which can offer sufficient guarantees for specific tasks or applications. In the context of multimodal representation learning, several recent studies have explored identifiability results [48, 13, 45], but largely neglect the presence of cross-modal misalignment. In contrast, our work explicitly models misalignment and adopts a block-identifiability definition to characterize the extent to which semantic factors can be recovered up to an invertible mapping.
@@
Invariant representation learning (IRL) seeks to learn representations that remain robust under distributional shifts between environments [2, 15, 18], particularly in settings where empirical risk minimization (ERM) [66] fails to generalize out of distribution. In the absence of such challenge, ERM is sufficient for in-distribution prediction, rendering the objective of IRL unnecessary. From a causal perspective, learning invariant representations, or more ambitiously, autonomous mechanisms [60], requires observing sufficient variations in the latent factors underlying data [41]. Such variability can be introduced through interventional data [39, 33], exchangeability assumptions [56], or the use of auxiliary variables such as domain indices [80, 42]. However, direct interventions on latent variables are typically infeasible in real-world observational data, and auxiliary variable methods often require access to a large number of diverse environments to ensure identifiability—an assumption that is often challenging to satisfy in practice. Our work sheds light on an alternative approach: by leveraging the inherent flexibility of text supervision, we show that manipulating biases, specifically through selective omission or semantic perturbation in text, can serve as a controllable proxy for environmental variation.
@@ C.1 Lemmas
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(f_{x},f_{t})=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\left[\|f_{x}(\mathbf{x})-f_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\right]-\frac{1}{2}\Bigl(H\bigl(f_{x}(\mathbf{x})\bigr)+H\bigl(f_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr)\Bigr),$ (6)
@@
$f_{x}^{*}=\mathbf{d}\circ(g^{-1}_{x})_{\mathbb{I}^{c}_{\rho}}:\mathcal{X}\to(0,1)^{n},$ (7)
$f_{t}^{*}=\mathbf{d}\circ(g^{-1}_{t^{\scriptscriptstyle(\theta)}})_{\mathbb{I}^{c}_{\rho}}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n},$ (8)
@@
  • $g_{x}$ and $g_{t^{\scriptscriptstyle(\theta)}}$ denote the true underlying generative mappings for images and paired text, respectively, as described in §˜3.
@@
We prove that the candidate functions $f_{x}^{*}$ and $f_{t}^{*}$ in Equations˜7 and 8 yield $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(f_{x}^{*},f_{t}^{*})=0$. Substituting these candidate functions into the loss in Eq.˜6, we have
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(f_{x}^{*},f_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\Bigl[\|f_{x}^{*}(\mathbf{x})-f_{t}^{*}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\Bigr]$
$\quad-\frac{1}{2}\Bigl(H\bigl(f_{x}^{*}(\mathbf{x})\bigr)+H\bigl(f_{t}^{*}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr)\Bigr).$
By the invertibility of the generative processes $g_{x}$ and $g_{t^{\scriptscriptstyle(\theta)}}$ (see §˜3), we may change variables to express the expectation over the latent variables:
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(f_{x}^{*},f_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\sim p_{\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}}}\Bigl[\|\mathbf{d}(\mathbf{s}_{\mathbb{I}^{c}_{\rho}})-\mathbf{d}(\tilde{\mathbf{s}}_{\mathbb{I}^{c}_{\rho}})\|_{2}\Bigr]$
@@
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(f_{x}^{*},f_{t}^{*})=0.$
Therefore, the global minimum of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ is achieved at 0 by the given function pairs $f_{x}^{*}$ and $f_{t}^{*}$, completing the proof. ∎
@@
Let $h:\mathcal{U}\to\mathcal{V}$ be a smooth map between simply connected, open $\mathcal{C}^{1}$ manifolds $\mathcal{U},\mathcal{V}\subseteq\mathbb{R}^{n}$. Suppose that $\mathbf{u}$ is a random variable taking values in $\mathcal{U}$ with a smooth probability density that is strictly positive a.e.. If the pushforward $\mathbf{v}=h(\mathbf{u})$ is uniformly distributed on $\mathcal{V}$, then $h$ is a global diffeomorphism; in other words, $h$ is bijective and depends on every component of $\mathbf{u}$. [^8] We do not claim originality for this result due to its fundamental nature in topology and measure theory; rather, we detail it here as a tool for our subsequent arguments.
@@
Let $p_{\mathbf{u}}:\mathcal{U}\to\mathbb{R}$ and $p_{\mathbf{v}}:\mathcal{V}\to\mathbb{R}$ denote the probability density functions of $\mathbf{u}$ and $\mathbf{v}$, respectively. Since $\mathbf{v}=h(\mathbf{u})$, the change-of-variables formula (refer to, e.g., [5]) yields
$p_{\mathbf{v}}(\mathbf{v})=p_{\mathbf{u}}(\mathbf{u})\cdot\Bigl|\det J(h)(\mathbf{u})\Bigr|^{-1},$
where $\mathbf{u}$ is any preimage of $\mathbf{v}$ under $h$. By assumption, $\mathbf{v}$ is uniformly distributed on $\mathcal{V}$; that is, there exists a constant $C>0$ such that
@@
Thus, for any $\mathbf{u}$ with $\mathbf{v}=h(\mathbf{u})$ we obtain
$\Bigl|\det J(h)(\mathbf{u})\Bigr|^{-1}=\frac{C}{p_{\mathbf{u}}(\mathbf{u})}.$
@@
$\Bigl|\det J(h)(\mathbf{u})\Bigr|^{-1}>0\quad\text{a.s.}\quad\forall\,\mathbf{u}\sim p_{\mathbf{u}},$
or equivalently, $\det J(h)(\mathbf{u})\neq 0$ a.e.. By the Inverse Function Theorem (see, e.g., [58]), this implies that $h$ is a local diffeomorphism.
Moreover, since $\mathcal{U}$ and $\mathcal{V}$ are simply connected, open $\mathcal{C}^{1}$ manifolds, standard covering space theory (refer to, e.g., the discussion around Theorem 1.38 in [21]) implies that $h$ is a covering map. The uniformity of $p_{\mathbf{v}}$ forces $h$ to be surjective (otherwise, some points in $\mathcal{V}$ would have zero density, contradicting uniformity). Since any covering map from a simply connected space is trivial, $h$ must be equivalent to the identity covering. In other words, $h$ is a homeomorphism onto $\mathcal{V}$ and hence both injective and surjective (i.e., a bijection).
Finally, the fact that the Jacobian determinant is nonzero a.e. guarantees that $h$ depends on all components of $\mathbf{u}$; if any component were omitted, the rank of the Jacobian would drop, contradicting non-singularity. Furthermore, by the Global Inverse Function Theorem (refer to, e.g., [59]), the inverse of $h$ is smooth.
In summary, $h$ is a global diffeomorphism from $\mathcal{U}$ onto $\mathcal{V}$. Consequently, it preserves all information of $\mathbf{u}$: every variation in $\mathbf{u}$ is reflected in $\mathbf{v}=h(\mathbf{u})$, and $\mathbf{u}$ can be uniquely and smoothly recovered from $\mathbf{v}$. This completes the proof. ∎
@@
  • First, we show that the objective function $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(f_{x},f_{t})$ achieves a global minimum value of 0. At this minimum, any pair of smooth functions $f_{x}$ and $f_{t}$ satisfying this condition must exhibit invariance across modalities. This invariance condition ensures that the learned image representations and text representations must align across all positive $\mathbf{x}$ and $\mathbf{t}^{\scriptscriptstyle(\theta)}$ pairs.
@@
Step 1 (Global minimum and invariance condition). Let $f_{x}:\mathcal{X}\to(0,1)^{n}$ and $f_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ be any smooth functions attaining the global minimum. Define the smooth mappings:
$h_{x}=f_{x}\circ g_{x},\quad h_{t}=f_{t}\circ g_{t^{\scriptscriptstyle(\theta)}}.$
@@
$\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x}}\,p_{\mathbf{t}^{\scriptscriptstyle(\theta)}|\mathbf{x}}}\left[\|f_{x}(\mathbf{x})-f_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\right]=\mathop{\mathbb{E}}_{(\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\sim p_{\mathbf{z}_{x}}p_{\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}|\mathbf{z}_{x}}}\left[\|h_{x}(\mathbf{z}_{x})-h_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\|_{2}\right]=0,$ (9)
$H\bigl(f_{x}(\mathbf{x})\bigr)=H\bigl(h_{x}(\mathbf{z}_{x})\bigr)=0,$ (10)
$H\bigl(f_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr)=H\bigl(h_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\bigr)=0.$ (11)